1. A body, weighing 5 kg, is displaced 30 m under the action of a force 30 N. parallel to the displacement. The coefficient of friction with the plane is 0.5. Find the work done by each force acting on the body, as well as the overall work.


1
Expert's answer
2021-04-08T08:51:38-0400
"W=Fs=(30)(30)=900\ N\\W_f=\mu mg=(0.5)(5)(9.8)=24.5\ N\\W_{total}=900-24.5=875.5\ N"